OpenStudy (anonymous):

0.6(y + 3) = 4.8 A. 5 B. 3 C. 13 D. 1.8

OpenStudy (anonymous):

rearrange the equation y + 3 = 4.8 / 0.6 = 8 y = 8 - 3 = 5

razor99 (razor99):

u have to slove for y right

razor99 (razor99):

yea emilie.li gave the correct anser

razor99 (razor99):

*answer

OpenStudy (anonymous):

thank you